Venture capital. Henry invested $12,000 in a new restaurant. When the restaurant was sold two years later, he received $27,000. Find his average annual return by solving the equation 12,000 (1 + r)2 = 27,000.

Answer by checkley77(12569)

12,000 (1 + r)^2 = 27,000
(1+r)^2=27,000/12,000
(1+r)^2=2.25
1+r=sqrt2.25
1+r=1.5
r=1.5-1
r=.5 or 50% ans.